# Physics-Informed Kolmogorov-Arnold Networks for multi-material elasticity problems in electronic packaging

### v1

@@ 2.3 Construction of admissible displacement fields
As described previously, DEM selects the displacement field that minimizes potential energy among all admissible displacement fields. The key to implementing admissible displacement fields lies in constructing displacement trial functions that naturally satisfy essential boundary conditions, embedding these constraints directly into the network architecture to ensure that the generated displacement field always satisfies essential boundary conditions during the optimization process.
Figure 1: DEM process overview.
As shown in Fig. 1, the general construction of admissible displacement fields can be expressed as
@@
where $\odot$ denotes element-wise product, $\mathbf{P}(\mathbf{x})$ is the boundary network that satisfies essential boundary conditions, and $\mathbf{F}(\mathbf{x};\boldsymbol{\theta})$ represents the neural network approximation. $\mathbf{D}(\mathbf{x})$ is the distance function, representing the shortest distance from the current point to the essential boundary $\Gamma_{u}$
@@
The key advantage of this formulation is that when coordinate points are located on the displacement boundary $\Gamma_{u}$, the distance function $\mathbf{D}(\mathbf{x})=0$. According to Eq. (13), the constructed displacement field output becomes exactly the prescribed essential boundary condition $\mathbf{P}(\mathbf{x})$. Consequently, essential boundary conditions are automatically satisfied during neural network training without requiring additional penalty terms and hyperparameter tuning. For detailed discussion of distance functions in this context, refer to [35].
$\mathbf{F}(\mathbf{x};\boldsymbol{\theta})$ is the generalized network, whose neural network parameters $\boldsymbol{\theta}$ are trained and optimized through the minimum potential energy loss function
@@
$=\underset{\boldsymbol{\theta}}{\arg\min}\left\{\int_{\Omega}\psi\left(\boldsymbol{\varepsilon}(\mathbf{u}^{\text{pred}}(\mathbf{x};\boldsymbol{\theta}))\right)\mathrm{d}\Omega-\int_{\Gamma_{\sigma}}\overline{\mathbf{p}}\cdot\mathbf{u}^{\text{pred}}(\mathbf{x};\boldsymbol{\theta})\mathrm{d}\Gamma\right\}.$
@@ 3 Methodology of PIKAN
Figure 2: Kolmogorov-Arnold Network schematic.
@@ 3.1 Kolmogorov-Arnold Networks (KANs)
In contrast, KANs move activation functions from nodes to edges, where each edge contains learnable univariate functions parameterized as B-spline functions. Instead of applying fixed activation functions at nodes after linear transformations (as in MLPs), KANs apply learnable activation functions on each edge connection. Nodes in KANs perform only summation operations on the outputs from incoming edges. This design enables each edge function to adaptively learn both the weight and activation behavior simultaneously, allowing for more flexible nonlinear transformations and enhanced model expressiveness and interpretability. During training, KAN updates the B-spline coefficients of these edge functions through backpropagation. The KAN architecture is illustrated in Fig. 2.
The core idea of the Kolmogorov-Arnold representation theorem is that any continuous multivariate function $f$ defined on a bounded domain can be represented as a combination of finitely many univariate continuous functions and additions
@@
where $\phi_{h,g}$ and $\Phi_{h}$ are univariate continuous functions. The starting point of KAN is to parameterize Eq. (16) explicitly, parameterizing all univariate functions $\phi_{h,g}$ and $\Phi_{h}$ as B-spline functions with learnable coefficients.
@@
To enhance the function approximation capability, scaling coefficients are introduced for each activation function [36]. For each activation function $\phi_{l,j,i}$, a corresponding scaling coefficient $m_{j,i}$ is defined. The scaling coefficient matrix $\mathbf{M}_{l}$ is
@@
Additionally, a linear transformation with weight matrix $\mathbf{W}_{l}$ and activation function $\sigma(\cdot)$ is introduced. The final output of the KAN layer is computed as
@@
In this work, when constructing KAN, we set the grid size $G_{j,i}$ and B-spline order $k_{j,i}$ to be consistent across all activation functions within each KAN layer. The trainable parameters in KAN layers are given in Table 1.
@@
$w_{j,i}$ Linear weights $N_{l}\times N_{l+1}$ Weights for auxiliary linear transformation with $\sigma(\cdot)$
@@
$\mathbf{x}^{(1)}$ $=\Phi_{0}^{*}(\mathbf{x}^{(0)})$
$\mathbf{x}^{(2)}$ $=\Phi_{1}^{*}(\mathbf{x}^{(1)})$ (26)
$\vdots$
$\mathbf{x}^{(L)}$ $=\Phi_{L-1}^{*}(\mathbf{x}^{(L-1)})$
@@ 3.2 Physics-Informed Kolmogorov-Arnold Networks (PIKANs)
PIKAN used in this work essentially implement the DEM using KAN network architecture, replacing the MLP traditionally employed in DEM with KAN. Consistent with traditional DEM applications in computational solid mechanics, the loss function construction is based on the principle of minimum potential energy. As shown in Eq. (15), KAN is trained by minimizing the energy-based loss function. Fig. 3 presents a schematic diagram of the complete PIKAN framework.
3.2.1 Geometric scale normalization
Since KAN uses B-spline functions to parameterize activation functions, it is necessary to preset the grid range of B-splines. In this work, the grid range is typically set to $[0,1]$ or [1,1][-1,1]. The algorithm takes the coordinates of sample points within the computational domain as inputs to the neural network. However, due to the arbitrary geometric dimensions of different models, the network inputs may exceed the preset B-spline grid range, necessitating input normalization to accommodate the grid range settings.
@@ 3.2.1 Geometric scale normalization
As shown in the upper left corner of Fig. 3, we consider a two-dimensional geometric model domain with characteristic length $L$ and characteristic height $H$, where $L\geq H$, and point $o$ is the coordinate origin. For any point $A(x_{A},y_{A})$ within the domain, we perform coordinate normalization
@@
The normalized coordinates $(\tilde{x}_{A},\tilde{y}_{A})$ are then used as inputs to the KAN network. This normalization process scales the input coordinates to ensure they are compatible with the preset B-spline grid range (e.g., $[0,1]$ or [1,1][-1,1]), allowing the B-spline activation functions to operate effectively within their defined domains.
We explain the necessity of geometric normalization through a mathematical analysis. For the model in Fig. 3, the left boundary is fixed ($u_{x}=u_{y}=0$). Without coordinate normalization, the admissible displacement field $u_{x}^{\text{pred}}(A)$ in the $x$-direction at point $A(x_{A},y_{A})$ is expressed as
$u_{x}^{\text{pred}}(A)=x_{A}\cdot\mathbf{F}\left(x_{A},y_{A};\boldsymbol{\theta}\right)$ (28)
During potential energy computation, displacements undergo differentiation with respect to coordinates. Differentiating Eq. (28) with respect to coordinate $x$ yields the normal strain in the $x$-direction
$\varepsilon_{xx}^{\text{pred}}(A)=\frac{\partial u_{x}^{\text{pred}}(A)}{\partial x}=\mathbf{F}\left(x_{A},y_{A};\boldsymbol{\theta}\right)+x_{A}\frac{\partial\mathbf{F}\left(x_{A},y_{A};\boldsymbol{\theta}\right)}{\partial x}$ (29)
The neural network output $\mathbf{F}\left(x_{A},y_{A};\boldsymbol{\theta}\right)$ is typically bounded within a finite range due to the activation functions used. However, the partial derivative term $\partial\mathbf{F}\left(x_{A},y_{A};\boldsymbol{\theta}\right)/\partial x$ can vary significantly with the input coordinates. When coordinate $x_{A}$ becomes large, the second term $x_{A}\frac{\partial\mathbf{F}}{\partial x}$ in Eq. (29) can become excessively large, causing the strain magnitude to be dominated by the coordinate value rather than the actual physical deformation. This leads to numerical instability and physically unrealistic results.
Introducing geometric normalization reformulates Eq. (28) as
$u_{x}^{\text{pred}}(A)=x_{A}\cdot\mathbf{F}\left(\tilde{x},\tilde{y};\boldsymbol{\theta}\right),$ (30)
where $\tilde{x}=x_{A}/L$ and $\tilde{y}=y_{A}/L$ are the normalized coordinates. The corresponding strain expression becomes
$\varepsilon_{xx}^{\text{pred}}(A)=\frac{\partial u_{x}^{\text{pred}}(A)}{\partial x}=\mathbf{F}\left(\tilde{x},\tilde{y};\boldsymbol{\theta}\right)+\tilde{x}\frac{\partial\mathbf{F}\left(\tilde{x},\tilde{y};\boldsymbol{\theta}\right)}{\partial x}.$ (31)
This normalization properly scales coordinate effects by ensuring that both terms in the strain expression remain within reasonable bounds. The coordinate multiplier $\tilde{x}(=x_{A}/L)$ in the second term is now normalized, preventing numerical instabilities and improving training convergence. Geometric normalization is therefore essential as it eliminates problematic scaling effects of coordinate derivatives, controls network inputs within a bounded range suitable for the B-spline grid, and enhances neural network fitting performance.
3.2.2 KAN layer output transformation
Since individual KAN layer outputs may exceed the preset B-spline grid range, potentially causing numerical issues in subsequent layers, we apply a $\tanh$ activation function to intermediate layer outputs to ensure they remain within the range [1,1][-1,1],
@@ 3.2.2 KAN layer output transformation
for $j=1,2,\ldots,N_{l+1}$ and layers $l=0,1,\ldots,L-2$. Note that the final layer (L1L-1) does not require the $\tanh$ activation function, since the output displacement field represents physical quantities with unbounded ranges that should not be artificially constrained.
[…]
@@ 3.3 Sample point distribution strategies
We employ various numerical integration schemes to approximate the energy function, including Monte Carlo, trapezoidal, and Simpson integration. For these schemes, sample points follow uniform distribution throughout the computational domain, maintaining consistent spacing to form a regular grid. Additional boundary points can be incorporated along curved boundaries to better capture domain geometry, as illustrated by the purple points in Figs. 5 and 5.
@@
Fig. 5 illustrates the triangular integration scheme, where purple points represent the original uniform grid, red points are triangle centroids serving as neural network training points, and the blue triangle area in Fig. 5(b) shows the control area for sample point $\mathbf{x}_{k}$. The mesh density should be chosen based on problem complexity.
@@ 4 Multi-material elasticity problems
Material property discontinuities at interfaces can cause difficulties in learning smooth displacement fields, leading to poor prediction accuracy without properly enforcing interface continuity conditions. Consequently, existing research predominantly employs domain decomposition approaches, where separate neural networks are assigned to different material domains according to material composition. Each neural network independently predicts the displacement field within its respective material domain. Regardless of whether strong-form or energy-form loss functions are used, interface continuity conditions must be incorporated into the loss function to ensure solution accuracy and physical consistency. Consequently, the more materials involved, the more interface penalty terms required in the loss function, significantly increasing training complexity and computational cost. For complex multi-material structures, this approach may result in convergence failure or non-smooth solutions.
Recent work has demonstrated KAN’s effectiveness in solving PDEs [28] and heterogeneous problems [36]. Based on spline interpolation theory, B-splines naturally exhibit piecewise polynomial characteristics with controllable smoothness properties. In multi-material elasticity problems, displacement fields remain continuous across material interfaces while stress fields (and consequently displacement gradients) may exhibit discontinuities due to material property changes. The piecewise nature of B-spline functions makes KAN theoretically well-suited for capturing such solution behaviors without explicit domain decomposition.
@@
Figure 6: PIKAN flowchart for multi-material elasticity problems.

### v2

@@ 2.2 Construction of admissible displacement fields
The DEM selects the displacement field that minimizes the potential energy among all admissible displacement fields. The key to realizing admissible displacement fields lies in constructing displacement trial functions that naturally satisfy essential boundary conditions and embedding this constraint directly into the network architecture to ensure that the displacement field consistently satisfies essential boundary conditions throughout the optimization process.
The general structure of an admissible displacement field can be expressed as
@@
where $\mathbf{P}(\mathbf{x})$ is the boundary function that satisfies essential boundary conditions. Note that for fixed boundaries, the prescribed essential boundary condition value is 0, i.e., $\mathbf{P}(\mathbf{x})=\mathbf{0}$. $\mathbf{F}(\mathbf{x};\boldsymbol{\theta})$ is the generalized network obtained through training by minimizing the potential energy loss function. $\odot$ denotes element-wise multiplication. $\mathbf{D}(\mathbf{x})$ is the distance function, representing the shortest distance from the current coordinate point to the essential boundary $\Gamma_{u}$
@@
The boundary and distance functions have flexible implementation forms; in addition to neural networks, fitting methods such as radial basis functions can also be employed Wang et al. [2022b]. The key advantage of this construction is that when the coordinate point is located on the boundary $\Gamma_{u}$, the distance function $\mathbf{D}(\mathbf{x})=\mathbf{0}$, and the admissible displacement field obtained from Eq. (3) naturally satisfies the enforced essential boundary conditions. In the validation examples presented in Section 5 of this study, given that the essential boundary geometries are relatively simple, $\mathbf{D}(\mathbf{x})$ consistently adopts the analytical distance function form Sukumar and Srivastava [2022], thereby avoiding additional errors introduced by neural network approximation while accurately enforcing essential boundary conditions. This method enables the admissible displacement field to naturally satisfy boundary constraints during generalized network training, without requiring the introduction of boundary condition penalty terms and associated hyperparameter tuning in its loss function. For detailed discussion on admissible displacement fields, please refer to Sukumar and Srivastava [2022].
The neural network parameters of the generalized network $\mathbf{F}(\mathbf{x};\boldsymbol{\theta})$ are trained and optimized by minimizing the potential energy loss function
@@
$=\underset{\boldsymbol{\theta}}{\arg\min}\left\{\int_{\Omega}\psi\left(\boldsymbol{\varepsilon}(\mathbf{u}^{\text{pred}}(\mathbf{x};\boldsymbol{\theta}))\right)\mathrm{d}\Omega-\int_{\Gamma_{\sigma}}\overline{\mathbf{p}}\cdot\mathbf{u}^{\text{pred}}(\mathbf{x};\boldsymbol{\theta})\mathrm{d}\Gamma\right\}$
@@ 3 Methodology of PIKAN
Figure 1: Kolmogorov-Arnold Network schematic (adapted and redrawn from the original KAN architecture Liu et al. [2024]).
@@ 3.1 Kolmogorov-Arnold Networks (KANs)
In contrast, KANs move activation functions from nodes to edges, where each edge contains learnable univariate functions parameterized as B-spline functions. Instead of applying fixed activation functions at nodes after linear transformations (as in MLPs), KANs apply learnable activation functions on each edge connection. Nodes in KANs perform only summation operations on the outputs from incoming edges. Their core advantage lies in adaptive learning capability for complex functional relationships and the resulting potentially higher parameter efficiency (i.e., achieving comparable accuracy with fewer parameters) Liu et al. [2024]. It must be objectively noted that choosing KANs over MLPs involves trade-offs. In many benchmark tests, their accuracies are often comparable Shukla et al. [2024]. The primary advantage of KANs lies in potentially mitigating spectral bias and more efficiently learning high-frequency features Meshir et al. [2025], Wang et al. [2025a], but the more complex gradient computations of B-spline functions during training increase computational cost Liu et al. [2024]. The piecewise function approximation characteristics of KANs are better suited for handling solution field discontinuities at multi-material interfaces, which traditional MLPs struggle to naturally capture. The KAN architecture is illustrated in Fig. 1.
The core idea of the Kolmogorov-Arnold Representation Theorem (KART) is that any continuous multivariate function f defined on a bounded domain can be expressed as a composition of finite univariate continuous functions and addition operations
@@
This provides theoretical inspiration for expressing multivariate functions through univariate function compositions. The design of KANs parameterizes the univariate functions $\phi$ and $\Phi$ in the theorem as learnable B-spline functions and extends them into nested, multi-layer network structures to enhance expressiveness. However, this deep KAN architecture goes beyond the scope of the original KART. As a new network paradigm, its theoretical validity is established and explored by the approximation theorem proposed and developed in the original paper itself Liu et al. [2024]. This work focuses on the application of this architecture, and the following will introduce its network structure and implementation details in detail.
@@
To enhance the function approximation capability, scaling coefficients are introduced for each activation function Wang et al. [2025b]. For each activation function $\phi_{l,j,i}$, a corresponding scaling coefficient $m_{j,i}$ is defined. The scaling coefficient matrix $\mathbf{M}_{l}$ is
@@
Additionally, a linear transformation matrix $\mathbf{W}_{l}$ and nonlinear activation function $\sigma(\cdot)$ are introduced to further enhance the network’s expressive and fitting capabilities. The final output of the KAN layer is computed as
@@
Following common practice to simplify implementation and training, this work adopts a uniform configuration where all activation functions within each KAN layer share the same grid size $G_{j,i}$ and B-spline order $k_{j,i}$. The trainable parameters in KAN layers are given in Table 1.
@@
$w_{j,i}$ Linear weights $N_{l}\times N_{l+1}$ weights for the nonlinear activation function $\sigma(\cdot)$
@@
$\mathbf{x}^{(L)}=\mathbf{\Phi}_{L-1}^{*}\left(\mathbf{\Phi}_{L-2}^{*}\left(\cdots\mathbf{\Phi}_{0}^{*}\left(\mathbf{x}^{(0)}\right)\cdots\right)\right)$ (16)
@@ 3.2 Physics-Informed Kolmogorov-Arnold Networks (PIKANs)
As discussed in the Introduction, PIKAN has emerged as a widely explored new paradigm in the field of physics-informed machine learning Liu et al. [2024], Shukla et al. [2024], Wang et al. [2025b], Jacob et al. [2025], Toscano et al. [2025a], Mostajeran and Faroughi [2025], Faroughi and Mostajeran [2025]. This section presents the specific PIKAN framework adopted in this study. As illustrated in Fig. 2, KAN replaces the MLP traditionally used in the Deep Energy Method for computational solid mechanics.
Since KANs employ B-spline activation functions defined on preset grids ([0,1] or [-1,1]), and the algorithm takes sample point coordinates within the computational domain as neural network inputs, while actual computational models have arbitrary geometric dimensions that may exceed the preset grid range, normalization of network inputs is necessary. Geometric normalization is crucial for ensuring training stability and convergence speed Toscano et al. [2025a], Mostajeran and Faroughi [2025], Wang et al. [2023], and its theoretical necessity has been thoroughly explored through rigorous analyses based on frameworks such as Neural Tangent Kernel (NTK), particularly in studies of Chebyshev KANs Mostajeran and Faroughi [2025], Faroughi and Mostajeran [2025]. This work implements this standardization strategy in specific problems based on these established theoretical foundations.
@@
As shown in the upper left corner of Fig. 2, we consider a 2D geometric domain with characteristic length $L$ and characteristic height $H$, where $L\geq H$, and point o is the coordinate origin. For any point $A(x_{A},y_{A})$ within the domain, we perform coordinate normalization
@@
The normalized coordinates serve as actual inputs to the KAN, ensuring B-spline activation functions operate effectively within their definition domain.
To mitigate training instability caused by KAN layer outputs potentially exceeding the definition domain of subsequent spline functions, this work adopts a validated approach of applying tanh activation functions to constrain intermediate layer outputs Shukla et al. [2024], Mostajeran and Faroughi [2025]
@@
for $j=1,2,\ldots,N_{l+1}$ and layers $l=0,1,\ldots,L-2$. Note that the final layer (L1L-1) does not require tanh activation, as the output displacement field represents physical quantities with inherently unbounded ranges that should not be artificially constrained.
@@ 3.3 Sample point distribution strategies
We employ various numerical integration schemes to approximate the energy function, including Monte Carlo, trapezoidal, and Simpson integration. For these schemes, sample points follow uniform distribution throughout the computational domain, maintaining consistent spacing to form a regular grid. Additional boundary points can be incorporated along curved boundaries to better capture domain geometry, as illustrated by the purple points in Figs. 4 and 4.
@@
Fig. 4 illustrates the triangular integration scheme, where purple points represent the original uniform grid, red points are triangle centroids serving as neural network training points, and the blue triangle area in Fig. 4(b) shows the control area $S_{k}$ for sample point $\mathbf{x}_{k}$. The mesh density should be chosen based on problem complexity.
@@
The choice of sampling strategy and numerical integration scheme directly affects loss function computation. For classical numerical integration schemes based on regular sample point distributions (Monte Carlo, trapezoidal, Simpson), this study employs their standard-defined weight coefficients to compute integral approximations. For Delaunay triangulation-based numerical integration, depending on different sample points (triangle centroids or uniformly distributed points), corresponding area coefficients are selected to approximate the energy integral. Section 5 systematically compares computational results across different schemes.
@@ 4 Multi-material elasticity problems
Recent work has demonstrated KAN’s effectiveness in solving PDEs Liu et al. [2024] and heterogeneous problems Wang et al. [2025b]. Based on spline interpolation theory, B-splines naturally exhibit piecewise polynomial characteristics with controllable smoothness properties. In multi-material elasticity problems, displacement fields remain continuous across material interfaces while stress fields (and consequently displacement gradients) may exhibit discontinuities due to material property changes. The piecewise nature of B-spline functions makes KAN theoretically well-suited for capturing such solution behaviors without explicit domain decomposition.